### v1

# Remote Sensing Image Intelligent Interpretation with the Language-Centered Perspective: Principles, Methods and Challenges

Abstract
The mainstream paradigm of remote sensing image interpretation has long been dominated by vision-centered models, which rely on visual features for semantic understanding. However, these models face inherent limitations in handling multi-modal reasoning, semantic abstraction, and interactive decision-making. While recent advances have introduced Large Language Models (LLMs) into remote sensing workflows, existing studies primarily focus on downstream applications, lacking a unified theoretical framework that explains the cognitive role of language. This review advocates a paradigm shift from vision-centered to language-centered remote sensing interpretation. Drawing inspiration from the Global Workspace Theory (GWT) of human cognition, We propose a language-centered framework for remote sensing interpretation that treats LLMs as the cognitive central hub integrating perceptual, task, knowledge and action spaces to enable unified understanding, reasoning, and decision-making. We first explore the potential of LLMs as the central cognitive component in remote sensing interpretation, and then summarize core technical challenges, including unified multimodal representation, knowledge association, and reasoning and decision-making. Furthermore, we construct a global workspace-driven interpretation mechanism and review how language-centered solutions address each challenge. Finally, we outline future research directions from four perspectives: adaptive alignment of multimodal data, task understanding under dynamic knowledge constraints, trustworthy reasoning, and autonomous interaction. This work aims to provide a conceptual foundation for the next generation of remote sensing interpretation systems and establish a roadmap toward cognition-driven intelligent geospatial analysis.
Index Terms:
Language-Centric Remote Sensing Interpretation, Global Workspace Theory, Multimodal Representation, Knowledge Association, Reasoning and Decision-Making.
I Introduction
Language and vision, as two fundamental paradigms of human cognition, have maintained a dialectical relationship throughout the development of civilization. Martin Jay[1] reveals the profound impact of the privileging of vision on modern cognitive science, where the visual system, due to its intuitive spatial perception, has been granted epistemological primacy within the framework of empiricist cognition.In contrast, the centrality of language is rooted in the Platonic tradition of idealism and underwent a modern transformation through the works of Descartes, Kant, and ultimately Wittgenstein’s ”linguistic turn”, which emphasized the abstract capacity of language as the foundation of rational cognition. The interplay between these two cognitive paradigms is also reflected in the history of biological evolution: the emergence of visual organs triggered the Cambrian explosion of biodiversity, while the birth of language marked the beginning of the human intelligence revolution. Kant, in his famous dictum ”Thoughts without content are empty, intuitions without concepts are blind.”[2] revealed the complementarity between the two: Vision offers the richness of concrete perception, while language provides the stability of abstract representation. Together, they form a closed loop through which humans cognitively engage with the physical world. In the interdisciplinary field of remote sensing science and artificial intelligence, this duality of cognitive paradigms is mirrored in the evolutionary trajectory of image interpretation technologies. Current mainstream intelligent remote sensing interpretation systems follow a vision-centered paradigm, relying on neural networks to extract pixel-level features and achieving remarkable progress in single-task scenarios such as object detection[3] and land-cover classification[4].
However, this paradigm faces three fundamental limitations. First, multimodal tasks—such as cross-modal retrieval and reasoning-based decision-making—often rely on the stacking of separate models, lacking a truly unified representational space. The root cause lies in the vision-centered paradigm’s dependence on spatiotemporal alignment across different modalities. For example, X Sun et al.[5] trained a multimodal model using jointly aligned optical, SAR, and nDSM data to ensure that information from different modalities could align and fuse within the same spatiotemporal reference frame. However, in most multimodal task scenarios, inherent mismatches across modalities—such as differences in acquisition methods, temporal synchronization, and spatial resolution—make it extremely challenging to construct a shared, unified representational space. Second, the interpretation process exhibits a unidirectional open-loop characteristic, lacking a dynamic feedback mechanism with the physical world. Due to the difficulty of effectively verifying the accuracy of results in an end-to-end interpretation approach, the model cannot adaptively adjust based on real-world feedback during the inference process. This lack of closed-loop validation mechanisms leads to cumulative errors when the model faces complex and variable real-world scenarios, impacting its robustness and generalization ability. Third, visual representations are constrained by the ”seeing is believing” perceptual limitations, making it difficult to extract deeper semantic relationships beyond overt features. CREPE[6] research indicates that visual models struggle to capture the implicit semantic relationships (such as functional attributes and conceptual links) between targets during recognition, limiting their abstract reasoning abilities. This limitation creates inherent deficiencies in visual models when it comes to understanding complex scenes, reasoning about causal relationships, or performing high-level cognitive tasks, thereby restricting their knowledge expression and reasoning capabilities in multimodal fusion. At its core, these limitations stem from the neglect of the conceptualization and structuring capabilities inherent in the language system, causing the interpretation results to remain at the surface level of vision and failing to form a closed-loop cognitive chain of ”Perception-Cognition-Decision”.
Figure 1: Overview of the organizational structure of review
With the advancement of large language model (LLM) technologies, research[7, 8] has shown that integrating vision and language models can help establish conceptual relationships among visual features. This insight has spurred exploration into remote sensing multimodal large language models (MLLMs), such as RS-LLaVA[9] and EarthGPT[10], which aim to combine visual information with language models to enhance the unification of task interfaces. These models leverage vision-language alignment techniques, enabling remote sensing image interpretation systems to effectively integrate multi-source data and support query and explanation through natural language interaction—thereby bridging the semantic understanding limitations of traditional methods. In recent years, remote sensing image interpretation has been gradually evolving toward multimodal reasoning models, such as EarthMarker[11], which aim to endow interpretation systems with stronger reasoning and decision-making capabilities. Existing review articles[12, 13, 14] have summarized this trend, focusing primarily on the technological development of MLLMs and their performance across various remote sensing interpretation tasks, highlighting improvements in accuracy and applicability. However, most of these reviews remain at the level of task-oriented application summaries and lack an in-depth analysis of the interaction mechanisms among different modalities, particularly the characteristics of multimodal interaction centered on the language modality.
Therefore, this review aims to fill this gap by systematically exploring the potential of a language-centered paradigm in remote sensing image interpretation. Inspired by the Global Workspace Theory[15] (GWT) of human cognitive models, we propose a theoretical framework for a global workspace-driven, cyclically interactive interpretation system. This framework demonstrates how language can be used to organize the interactions among interpretation tasks, sensors, and models, thereby achieving a cognitive closed loop that overcomes the limitations of current interpretation systems. It identifies three major challenges faced by current technologies in constructing a language-centered intelligent interpretation framework: the bottleneck of unified multimodal representation, the difficulty of multi-source knowledge association, and the limited capabilities in reasoning and decision-making. This work provides a theoretical foundation for building language-centered remote sensing interpretation systems and seeks to chart a path for multimodal remote sensing intelligence that ”returns to the essence of cognition.” The main contributions of this paper are as follows:
  1. 1.
    Rethinking Remote Sensing Image Interpretation from a Cognitive Perspective: Starting from the human cognitive process, this work analyzes the cognitive limitations of current remote sensing image interpretation technologies and emphasizes a language-centered ”Perception–Cognition–Decision” paradigm, laying a theoretical foundation for future research.
  2. 2.
    Global Workspace-Driven Interpretation Framework: In light of recent advances in LLMs, this work explores the potential application of GWT in remote sensing image interpretation. It provides a systematic review of relevant technologies and analyzes the feasibility of implementing such a framework within the current technological landscape.
  3. 3.
    Key Challenges and Technical Bottlenecks in Building a Language-Centered Intelligent Interpretation Framework: This work summarizes the core issues faced in implementing language-centered remote sensing interpretation frameworks. It focuses on three major challenges:
    • The difficulty of constructing unified multimodal representations;
    • The lack of effective mechanisms for associating multi-source knowledge;
    • The limited reasoning and decision-making capabilities.
    Potential research directions to address these challenges are also discussed.
  4. 4.
    Outlook and Emerging Research Opportunities: Finally, this review summarizes and discusses the new opportunities brought by the paradigm shift toward language-centered interpretation frameworks in the field of remote sensing image understanding. It also proposes potential research directions, including novel areas not covered by previous surveys.
The overall structure of this review is illustrated in the Fig.1.
II Language-Centered
Remote Sensing Intelligence Interpretation
II-A Global Workspace-Driven Interpretation Mechanism
Human cognition of the physical world is a cyclical and interactive process, with bottom-up (perception forming cognition) and top-down (cognition guiding perception) pathways being the two fundamental ways humans understand the world. These processes are unified through consciousness. How consciousness bridges the physical world and human cognition through these two pathways is explained by various theories. The GWT is one of the mainstream perspectives, proposed in 1998 by Baars[15]. It posits that there exists a ”Global Workspace” in the brain as a shared platform for gathering and interacting information, capable of integrating information from different sources to support complex cognitive functions. Within this framework, bottom-up sensory information and memories in the brain converge and associate in the global workspace, where reasoning and further action decisions are made, guided by human intention. Ultimately, the brain sends out action commands that control the body’s interaction with the external world, enabling a response to the complex environment. For example, when a person is driving a car to find a parking space, first, visual perception detects road signs, parking spaces, and other vehicles’ positions, auditory perception senses traffic noise, and tactile perception via the steering wheel provides feedback on the road surface. These sensory inputs are transmitted to the brain via the nervous system in a bottom-up manner and integrated in the global workspace. In the global workspace, reasoning driven by human intention (to find an available parking space) makes decisions based on existing knowledge (surrounding environment information, current vehicle position), determining the next actions (steering the wheel, pressing the brake/accelerator, etc.). This action leads to environmental changes, which in turn provide the brain with new sensory information, creating a continuous cycle of interaction.
Figure 2: Correspondence Between Human Cognition and Intelligent Remote Sensing Interpretation
As shown in Fig.2, in the human cognitive process, multi-source information from complex physical scenes is captured by sensory organs (such as the eyes, ears, skin, etc.) and transmitted to the human cognitive global workspace. The global workspace generates intervention intentions based on this sensory information and interacts with the physical world through effectors (such as hands and feet). In this interaction, the physical space and cognitive space influence each other, continually driving human thinking and learning processes. Throughout this process, humans continuously refine their skills, gradually improving their ability to handle complex tasks.
Similarly, remote sensing intelligence interpretation of the physical world is also a cyclical and interactive process. Temporal and spatial dynamics, along with the perception of complex physical scenes that are constantly changing, are captured by remote sensing observation systems and quantified into digital representations. After analyzing and modeling these data, the information from the physical scene can be transmitted to the global workspace of remote sensing intelligence interpretation. Under the decision support of the global workspace, the system issues control commands to the remote sensing observation devices, actively acquiring more sensory data based on the needs of the interpretation task, thereby continuously adjusting and optimizing the interpretation of the physical scene.
In general, the remote sensing observation system transforms physical information from the real world into digital representations through a bottom-up perception process, transmitting it to the interpretation space for processing. When these sensory inputs are insufficient to meet the demands of the interpretation task, the remote sensing intelligent interpretation model needs to adjust the observation strategy top-down based on the current interpretation requirements to further acquire sensory data. For example, in a ship target interpretation task, the remote sensing intelligent interpretation system captures images of ships on the sea surface through sensors mounted on drones and inputs the digitalized representations into the system. The intelligent interpretation system analyzes these data and, combined with task knowledge, captures key features (such as the hull number) for identifying the ship type. However, when the hull number cannot be clearly obtained due to occlusion from the viewpoint, the system directs the drone to interact with the physical world by moving, adjusting the focal length, or switching sensory modalities, performing targeted actions to capture key details of the hull number.
Figure 3: Cyclic Interactive Interpretation Driven by a Global Workspace
The cyclic and interactive interpretation process includes a bottom-up remote sensing image acquisition procedure and a top-down reasoning and decision-making process guided by the interpretation task. This process can be driven by a Global Workspace and formally represented as Cyclic Interaction Interpretation:
$C=\left(G(P,K,T),\ A\right)$
The entire process can be abstracted as shown in Fig.3, where:
  • $P$ representing the perception space, denotes information such as scene context, object characteristics, and spatiotemporal attributes. This space captures a detailed description of geo-objects within the observation environment, including their spectral signatures, geometric properties, texture features, and temporal dynamics.
  • $K$ represents the knowledge space, referring to supplementary information derived from remote sensing expert knowledge. This includes prior knowledge of object features, geographical distribution patterns, seasonal variation rules, and more. Such knowledge can be acquired through expert experience or accumulated historical data, and is used during the interpretation process to enhance the understanding and reasoning over observed information.
  • $T$ denotes the task space, representing the user’s interpretation requirements—i.e., the specific objectives the interpretation system is expected to accomplish. Tasks within the task space may range from simple ones (such as identifying the object category in an image) to more complex ones (such as recognizing the fine-grained category of a ship in the image when key visual features are missing).
  • AA represents the action space, referring to the set of executable actions available to the sensing system. These actions include adjustments to the sensor’s resolution, intrinsic parameters, sensing modalities, combinations of modalities, as well as changes in observation angles and orientations.
  • $G$ denotes the Global Workspace, which integrates information from $P$ (Perception Space), $K$ (Knowledge Space), and $T$ (Task Space), and outputs control instructions to $A$ (Action Space).
In the cyclic interaction interpretation process, the Global Workspace encodes and integrates the visual representations from the Perception Space, the remote sensing expert knowledge from the Knowledge Space, and the interpretation requirements from the Task Space into a unified representation. It then establishes knowledge associations between perceptual information and expert knowledge. Guided by the interpretation requirements, Global Workspace performs reasoning and decision-making based on the perceptual information and expert knowledge, thereby controlling subsequent interactive actions.
II-B The Potential of Applying LLMs as the Global Workspace
The global workspace drives the cyclic interaction interpretation process through its capabilities in unified encoding, knowledge association, and reasoning-based decision-making. With the rapid development of LLMs, such as LLaMA[16], GPT[17] and DeepSeek[18], these models have been shown to effectively encode world knowledge and exhibit logical reasoning capabilities. Meanwhile, the integration of LLMs has also led to significant advancements in areas such as unified multimodal modeling, knowledge base construction, and embodied intelligence.
Therefore, this paper argues that LLMs possess strong potential to serve as the global workspace within remote sensing intelligent interpretation systems, for the following reasons:
LLMs possess the ability to construct unified representations across multimodal data
Different modalities essentially express the same entity from distinct perspectives. Based on this view, the Platonic Representation Hypothesis suggests that representations across modalities tend to converge consistently, and using language as the central convergence point facilitates cross-modal information association and reasoning.In the visual domain, pioneering vision-language models such as CLIP[7] and ALIGN[19] jointly train vision and language modalities by mapping both into a shared semantic space. This significantly enhances the model’s ability to understand deep semantic associations between visual elements, confirming that although image and text are distinct in modality, they can converge through shared semantic concepts.BLIP[20] further supports this by generating descriptive textual captions for visual inputs, thereby demonstrating that visual information can be effectively described through language and enabling the construction of a shared vision-language representation space.In the broader multimodal context, LanguageBind[21] employs contrastive learning to align a variety of modalities—including video, infrared, depth, and audio—with language, enabling their embedding into a unified language-centered space. Furthermore, AllSpark[22] successfully aligns representations from 13 different modalities into the language modality.These works collectively validate: (1).That heterogeneous modality data can be embedded into a unified space via the language modality; (2).That diverse modality representation spaces can be integrated through a language-centered framework.
LLMs possess the ability to associate perceptual information with existing knowledge
Through training on large-scale corpora, LLMs encode the logical relationships within language, manifesting in their strong logical reasoning capabilities[23, 24]. This enables LLMs to first comprehend the perceptual information in multimodal data and then match it with previously learned knowledge, thus driving advancements across multiple fields. In terms of information understanding, LLMs can not only understand tasks in algebra, calculus, and geometry by recognizing objectives and variable conditions, thereby constructing appropriate problem-solving logic, but also handle geographic data mining and analysis tasks, gaining deep insights into task objectives such as spatial patterns, geographic entity relationships, and trend analysis.Regarding knowledge association, the knowledge linking ability of LLMs stems from the vast world knowledge, linguistic patterns, and conceptual associations accumulated during their training process. This knowledge serves as the model’s ”long-term memory”, allowing it to quickly integrate relevant information and provide explanations when faced with complex inputs. The PaLM model proposed by Chowdhery et al.[25] demonstrates that the large-scale parameters of LLMs are crucial for enhancing retrieval capabilities. This allows the model to associate learned knowledge with minimal prompting in complex contexts, enabling it to make accurate responses.
LLMs possess the ability to handle complex task reasoning and decision-making
Reasoning is essentially the process of drawing conclusions from known information or premises through logical inference chains. LLMs achieve reasoning and decision-making for complex tasks by constructing inference chains and progressively deducing from the associated information.Research by Wei et al.[26] explored the role of chain-of-thought prompting in large language models, demonstrating that by clearly defining reasoning links, LLMs can not only better understand the task itself but also efficiently solve complex tasks that require multiple reasoning steps. Moreover, reasoning and decision-making for complex tasks often require LLMs to dynamically adjust their reasoning chains based on task objectives. Relevant studies have shown that LLMs can optimize their decision-making process by adjusting their reasoning chains.An example of this capability is the DeepSeek-R1[27] model which can adjust its reasoning chain and refine its reasoning strategy before responding. By incorporating feedback information, it can correct previous errors, thereby solving complex tasks. For instance, during the AIME 2024 test, DeepSeek-R1[27] successfully solved 79.8% of high-difficulty problems by reflecting on and adjusting its reasoning steps.
Figure 4: Paradigm Shift in Remote Sensing Image Interpretation: From Vision-Centered to Language-Centered
II-C Redefining Intelligent Interpretation of Remote Sensing Imagery
Currently, the mainstream approach to intelligent interpretation of remote sensing imagery follows a vision-centered passive perception paradigm, typically employing an end-to-end inference mechanism: input data is processed by a model, which then produces interpretation results based on pre-trained parameters. Under this paradigm, the model is usually tightly coupled with the specific task, meaning that an independent model must be trained for each individual task, and that model relies heavily on data from a particular application scenario. This highly specialized design significantly limits the model’s generalization ability. Especially when dealing with complex tasks involving multiple decision-making processes, it often becomes necessary to design and train separate models for each decision flow to meet varying task requirements. As a result, the overall system construction cost increases substantially, and the flexibility to adapt to changing scenarios is greatly reduced.
In contrast, bottom-up (perception-driven cognition) and top-down (cognition-guided perception) processes represent two fundamental pathways through which humans understand the world. Combining these two paths enables the formation of a ”Perception + Cognition” cyclic framework. This framework not only allows for the in-depth extraction of visual information at different levels of granularity, but also supports the construction of a multi-task reasoning chain by associating perceptual input with existing knowledge, thereby enabling complex visual reasoning tasks.
As illustrated in Fig.4, traditional ship target recognition methods rely on a single visual model and follow an end-to-end production pipeline, where task forms are fixed. In contrast, a language-centered approach to intelligent interpretation of remote sensing imagery allows users to interact with the model through natural language and actively assess the progress of interpretation tasks. When key information is missing, the model can interact with sensors to acquire additional perceptual data, thereby significantly enhancing the flexibility and adaptability of the interpretation system.
Therefore, this paper redefines the process of intelligent interpretation of remote sensing imagery as follows: it is an interpretation process centered on language-organized interaction among tasks, sensors, and models, and driven by cyclic, multi-round dialog-based reasoning. This process integrates perception and cognition to support flexible task configurations. Remote sensing image interpretation should no longer be centered on vision-driven passive perception, but should instead be realized through a multi-level framework that combines perception and cognition.
This framework effectively integrates perception-driven visual processing with cognition-driven semantic reasoning, enabling the interpretation system to go beyond basic image analysis and leverage external knowledge for task reasoning and information completion. Such a ”perception + cognition” interpretation framework not only has the potential to enhance the generalization capabilities of interpretation models but also provides an effective solution for addressing complex interpretation tasks.
II-D Advantages of Language-Centered Intelligent Interpretation of Remote Sensing Imagery
The language-centered intelligent interpretation paradigm emulates the fundamental way in which humans understand the world—by integrating the perception space, knowledge space, task space, and action space into a cyclic interpretation model capable of effective interaction with the physical world. In this paradigm, language is not merely a tool for information transmission, but serves as the central hub for knowledge integration, coordination, and decision-making. Through a language-driven cyclic interpretation process, the model can handle complex interpretation tasks in open environments, enabling a shift from passive environmental perception to active environmental interaction in remote sensing image interpretation.
The rapid advancement of LLMs makes this paradigm increasingly feasible. With their abilities in unified representation, knowledge association, and reasoning and decision-making, LLMs overcome the inherent limitations of vision-based models—limitations stemming from the tightly coupled nature of data, models, and tasks in traditional training pipelines. As a result, language-centered models demonstrate the following advantages in intelligent remote sensing image interpretation:
Language-centered interpretation models offer advantages in integrating multimodal perceptual data
Due to differences in their intrinsic mechanisms (e.g., spatial resolution, spectral characteristics) and extrinsic factors (e.g., temporal and spatial coverage), different modalities are often highly heterogeneous at the structural level, posing significant challenges for alignment and collaborative interpretation. Numerous studies have found that in human cognition, information from various modalities—such as auditory, tactile, olfactory, and visual—is ultimately conceptualized and represented through language. Language serves not only as a carrier of perception and understanding but also as the medium for association, reasoning, and expression. This suggests that language not only encodes our perception of the world with precision but also implicitly establishes associations across different modalities. Although remote sensing data types such as optical, infrared, hyperspectral, SAR, and LiDAR differ in both intrinsic mechanisms and acquisition conditions, they can all be considered semantically as different views of the same geo-object. Recent multimodal language model research—such as CLIP[7], LanguageBind[21], and AllSpark[22] has validated the feasibility of embedding diverse modalities into a shared language space. Therefore, language-centered interpretation models possess the potential and advantages to align multimodal perceptual data semantically via the language modality and to effectively fuse heterogeneous information across modalities.
Language-centered interpretation models possess the ability to learn in open environments, enabling them to quickly adapt and integrate new information when encountering unfamiliar scenes or data
In contrast, traditional vision-based models primarily rely on supervised learning through the mapping between massive amounts of remote sensing imagery and a limited set of semantic annotations. This approach suffers from inherent theoretical limitations, as a finite set of semantic labels cannot provide sufficient information to thoroughly describe geo-objects and their relationships—thus restricting the model’s understanding of the world to the boundaries defined by the labeled knowledge. As a result, remote sensing vision models often struggle to perform effectively when faced with unlabeled or novel scenarios, making it difficult for them to generalize beyond predefined categories. In stark contrast, LLMs encode human knowledge that is inherently expressed in language. By associating perceptual information with a vast and structured knowledge base, LLMs can overcome the constraints imposed by limited annotations and significantly expand the range of learnable knowledge. For example, in ship recognition tasks, even when encountering previously unseen ship categories, a language-centered model can identify key component features from the image and reason about the specific type of vessel by associating those features with relevant prior knowledge. Specifically, different classes of ships, such as aircraft carriers, typically exhibit stable and class-distinguishing features. For instance, the number of ship islands can be used to distinguish the Queen Elizabeth-class aircraft carrier from other carriers. This is made possible because language enables compositional reasoning, allowing the model to associate various component features and derive logical conclusions.
Language-centered interpretation models offer clear advantages in handling complex interpretation tasks
Leveraging the conceptual relationships encoded within LLMs, such models can decompose complex tasks into simpler subtasks and perform reasoning and decision-making by integrating and associating diverse information within the language space. In contrast, vision-centered interpretation models typically require a separate, dedicated model for each complex task. This not only increases the complexity of model development and maintenance but also limits each model’s applicability to a narrow task scope, making it difficult to adapt to rapidly changing interpretation demands. For example, in remote sensing monitoring of flood disasters, an LLM can first analyze low-resolution remote sensing images collected by various sensors to identify the affected flood zones. It can then match these zones with spatial distribution maps of different functional areas—such as agricultural or industrial zones—to infer the degree of damage in each area. Based on this preliminary damage assessment, the LLM can dynamically mobilize relevant sensing resources, such as high-resolution imagery sensors, meteorological sensors, and geological sensors, to acquire more detailed and real-time remote sensing data for further analysis of the flood’s impact on each region. In parallel, by combining historical data with real-time sensor inputs, the model can assess the potential risk of secondary disasters, such as landslides or mudflows. Through this cyclic, step-by-step reasoning and decision-making process, the model can generate comprehensive predictions of both the flood’s impact and the likelihood of associated secondary hazards.
Language-centered interpretation models have the advantage of actively acquiring perceptual information
By deeply understanding the interpretation task, these models can assess whether the available perceptual information is sufficient and proactively seek additional data when needed. This capability stems from the model’s mechanisms for knowledge association and reasoning-based decision-making, allowing it to evaluate the relevance and completeness of information during interpretation and expand its knowledge scope when necessary to meet the demands of complex tasks. In contrast, traditional vision-based models are limited to the perceptual information contained within the imagery itself. When such information is insufficient or when complex scenarios require external knowledge support, vision models often struggle to adapt. For instance, in the task of observing and analyzing ship types, conventional imagery may fail to provide sufficient recognizable features due to weather conditions (e.g., clouds or fog) or suboptimal observation angles. In such cases, a language model can analyze the perceptual input—such as blurred images or missing hull numbers—and reason about the next best observation action to fulfill the interpretation goal. The model can then guide the system to perform more targeted sensing operations, such as issuing a command for a drone to switch to a high-resolution optical camera and move closer to the target for clearer imaging. Through this process, the language model can actively acquire critical information, enabling more effective completion of the interpretation task.
In summary, compared to traditional vision-centered interpretation models, language-centered models demonstrate clear advantages. Leveraging their rich knowledge association capabilities and powerful reasoning abilities, language models can actively acquire and integrate multi-source data, effectively addressing the challenges posed by open remote sensing scenarios and complex interpretation tasks. The shift from a vision-centered to a language-centered paradigm broadens the scope of intelligent remote sensing image interpretation and enhances the model’s adaptability and interpretive accuracy in contexts characterized by multi-source data, open environments, and complex task demands.
III A Language-Centered Framework for Intelligent Interpretation of Remote Sensing Imagery
Although the language-centered framework for intelligent interpretation of remote sensing imagery demonstrates significant advantages over traditional vision-centered approaches—particularly in handling complex and open remote sensing scenarios with greater flexibility, interactivity, and reasoning capability—it also faces a series of challenges in practical implementation, as illustrated in Fig.5. First, remote sensing imagery in the perception space encompasses multiple modalities, and the differences in spatial and temporal resolution, coverage, and spectral characteristics across these modalities make it difficult to unify their representations within a global workspace. Second, the perceptual information from the perception space, the interpretation tasks from the task space, and the internal and external prior knowledge in the knowledge space are often difficult to effectively associate. Finally, the knowledge elements themselves often exhibit complex logical relationships, making it challenging to perform accurate reasoning and to generate reliable decisions for subsequent actions.
Figure 5: Challenges and key problems in building a language-centered framework for intelligent remote sensing image interpretation.
In summary, intelligent interpretation of remote sensing imagery faces significant challenges in three key areas: multimodal unification, knowledge association, and reasoning and decision-making. Multimodal unification concerns the harmonization of data representations across different modalities and the alignment of their semantic meanings. The core of knowledge association lies in accurately understanding the interpretation task and retrieving relevant knowledge with precision. Reasoning and decision-making focus on dynamically analyzing interpretation progress based on perceptual inputs and associated knowledge, and on effectively planning observation strategies to optimize the interpretation process. This section will first provide an in-depth analysis of these core challenges, followed by a review of the current research status and recent advancements in each of these subdomains.
III-A Overview of Core Technical Challenges
III-A1 Unified Multimodal Representation
In a language-centered remote sensing intelligent interpretation system, the alignment of multimodal data representations is key to achieving effective cross-modal information fusion. The central challenge lies in how to map heterogeneous information from different sensing modalities into a unified global workspace for integration and analysis. However, aligning multimodal remote sensing data to a shared language space still faces several significant challenges:
Structural alignment across modalities
Non-linguistic modalities differ significantly from language modalities in terms of data representation formats, making it difficult to embed them into a shared semantic space. For example, RGB, MSI, and SAR imagery are typically represented as dense 2D matrices, while LiDAR point cloud data is stored as sparse 3D point sets. These structural inconsistencies can lead to loss of perceptual information when non-linguistic modalities are mapped into a 1D language-based semantic space, thereby affecting the effectiveness of multimodal fusion. Therefore, one of the fundamental challenges for language-centered unified representation is how to standardize the representation of non-linguistic modalities in a way that preserves their perceptual content while aligning them with language-based representations.
Semantic alignment across modalities
Language modalities carry human abstract conceptual systems, while non-language modalities are direct samplings of the physical world. This fundamental difference means that aligning their semantics typically requires additional human prior knowledge. However, in the field of remote sensing, non-language modality data is vast and rich in spatiotemporal information, making it extremely costly and labor-intensive to build precise alignment between language and non-language modalities through manual annotation. Furthermore, there are often significant mismatches in granularity and informational content between the two modalities. For instance, the term ”ship” in natural language typically refers to the entire vessel, whereas in remote sensing imagery, the representation of a ship may consist of various individual visual components. Direct cross-modal alignment without proper abstraction may result in semantic ambiguity. Thus, a core challenge for language-centered multimodal fusion is how to establish accurate semantic alignment mechanisms while minimizing reliance on manual annotation.
III-A2 Knowledge Association
The global workspace contains a complex composition of knowledge, and the ability to associate and integrate knowledge from different sources plays a critical role in solving complex remote sensing image interpretation tasks. Knowledge association is essential for enabling the system to make correct reasoning and decisions. Therefore, achieving accurate knowledge association is one of the primary challenges in building a language-centered intelligent interpretation system for remote sensing imagery. Knowledge association is a complex process that requires establishing effective links among perceptual information, task requirements, and both internal and external knowledge sources. Specifically, it involves two key steps: task understanding and knowledge retrieval.
Understanding remote sensing interpretation tasks
Task understanding is the prerequisite for accurate knowledge association. A remote sensing interpretation system must deeply comprehend the task requirements and relate them to perceptual information to determine which pieces of information are most relevant. For example, in a fine-grained ship classification task, different types of ships within the same general category often differ only in specific subtle features. Therefore, the interpretation system must leverage internal knowledge to first identify the coarse category of the target ship, and then analyze the image content to determine which visual features—such as the position or number of ship islands, or the hull number—are most helpful for fine-grained classification. This enables the system to more efficiently guide the execution of subsequent recognition steps.
Knowledge retrieval in remote sensing image interpretation
Knowledge retrieval is key to accurately extracting relevant information. Remote sensing interpretation systems must not only utilize the knowledge encoded within the model itself but also possess the ability to extract task-relevant knowledge from external multimodal sources. Continuing with the example of fine-grained ship classification, once the system has fully understood the task requirements, it must retrieve and extract information related to ship subcategories and their distinctive features from both internal and external knowledge sources. For instance, the Queen Elizabeth-class aircraft carrier is equipped with two ship islands—this type of feature can serve as crucial evidence in subsequent reasoning and decision-making processes and directly impacts the system’s interpretation performance. In practical terms, knowledge retrieval involves two core technical challenges: efficient knowledge encoding and reliable knowledge retrieval. The former requires the model to embed various forms of data—such as structured data, unstructured text, or temporal information—into its internal parameters, while ensuring that the information is well-organized and meaningfully associated to avoid knowledge confusion. The latter demands that the model be able to rapidly locate and retrieve information highly relevant to the current task from a vast internal and external knowledge space, by leveraging the complex associative relationships among knowledge elements.
III-A3 Reasoning and Decision-Making
Accurate and efficient reasoning and decision-making are key to driving the cyclic interactive interpretation process. Within the global workspace, reasoning and decision-making require the interpretation system to integrate task-relevant internal knowledge during the perception process. The system must first assess whether the currently available information is sufficient to accomplish the interpretation task, and then identify any missing information and formulate an optimal next-step sensing strategy. Through this process, the interpretation system can proactively select regions or targets that require further observation, thereby improving both sensing efficiency and interpretation accuracy.
Assessing the gap between external information and task objectives
One of the primary challenges faced by interpretation systems is how to evaluate the gap between the available external information and the requirements of the task. This gap assessment largely determines the system’s cognitive capability and perceptual efficiency. Only when the system can accurately assess the relevance and adequacy of the current external information can it make reasonable decisions for subsequent planning. The process of evaluating this gap is essentially a multi-level and complex computation. It requires both a precise understanding of the task objectives and a comprehensive analysis of the characteristics and variations in the external data. If this evaluation is inaccurate, the system’s reasoning path may deviate from the correct decision-making direction, leading to biased interpretation results and ultimately compromising task performance. Therefore, designing an efficient and accurate mechanism for information gap assessment—one that enables the system to maintain stable and precise judgment in complex and dynamic information environments—is a crucial prerequisite for achieving active perception.
Reasoning path optimization and action decision-making
A central challenge lies in designing an efficient decision-making mechanism that ensures the reasoning path leads to optimal perceptual efficiency. The purpose of the reasoning path is to progressively narrow the gap between external information and task objectives, allowing the system to dynamically adapt to environmental changes as it moves closer to its goal. However, in real-world physical environments, external conditions are often highly dynamic and influenced by numerous factors—such as the movement of target objects, changes in lighting, or variations in weather conditions within the sensing range. These variations cause perceptual information to continuously change in both space and time. In such contexts, the interpretation system cannot rely on fixed, pre-defined task strategies for information acquisition. Therefore, the system must possess flexible, adaptive capabilities that allow it to adjust its reasoning path in real time based on environmental changes, ensuring that it can still capture critical information effectively in dynamic conditions. As such, designing an adaptive decision-making mechanism that meets the practical demands of real-world tasks and guides the system to dynamically plan its reasoning and action path is another core research issue for achieving efficient reasoning and decision-making.
III-B Language-Centered Unified Multimodal Representation
In language-centered remote sensing intelligent interpretation systems, the key to unified multimodal representation lies in the effective mapping of heterogeneous information from different modalities into the language space. This requires not only the unification of data representation formats across modalities but also the accurate alignment of the semantics conveyed by each modality.
Therefore, the process of language-centered unified multimodal representation involves two core problems:
  • How to align the structural of different modalities?
  • How to align the semantics across modalities?
The structure of this section is illustrated in Fig.6, and representative methods corresponding to each component are summarized in Table I.
Figure 6: Structural organization of language-centered multimodal representation unification.
TABLE I: Unified multimodal representation structure and representative methods.
Content structure Category Representative algorithms
Structural Alignment Static Sequential Embedding VisualBERT[28], LXMERT[29], UNITER[30], RSGPT[31], GeoChat[32], RS-LLaVA[9]
Dynamic Hybird Embedding Oscar[33], VIMA[34], PaLM-E[35], TEOChat[36], SkySenseGPT[37]
Semantic Alignment Scene-Caption CLIP[7], UniCL[38], K-LITE[39], LaCLIP[40], SimVLM[41], BLIP[20], Qu et al.[42], RSICD[43], RSITMD[44], RemoteCLIP[45], GeoCLIP[46]
Region-Phrase ViLBert[47], RoIAlign[48], VL-BERT[49], ViLT[50], FGVP[51], Alpha-CLIP[52], ASM[53], Zou and Shi[54], GeoVG[55], RSVG[56], TEMO[57]
Scene-Instruction LLaVA[58], InstructBLIP[59], CAT[60], SEAL[61], RSVQA[62], VQA-TextRS[63], Prompt-RSVQA[64]
Implicit alignment Maniparambil et al.[65], LM4VisualEncoding[66], LAMP[67], GRAFT[68]
III-B1 Structural Alignment Across Modalities
Data from different modalities often exhibit significant differences in spatial resolution and temporal coverage. For example, the spatial resolution of remote sensing imagery may vary depending on the sensor type (e.g., RGB, MSI, SAR), while temporal coverage is influenced by the frequency of data acquisition. Traditional approaches typically represent each modality in distinct formats, making cross-modal fusion difficult and lacking effective channels for information interaction. However, with the success of Transformers[69] in natural language processing, an increasing number of studies have demonstrated that it is possible to achieve unified representation of multimodal data by constructing modality-specific embeddings. In this approach, the Transformer processes data from different modalities at the embedding layer, effectively mapping them—despite spatial and temporal differences—into a shared semantic space, thus enabling seamless fusion and collaboration across modalities. Depending on how the embeddings from different modalities are combined, this strategy can be categorized into two main types: Static Sequential Embedding, and Dynamic Hybrid Embedding.
Static Sequential Embedding
As illustrated in Fig.7(a), the static sequential embedding approach combines embeddings from different modalities in a fixed order and length after passing each modality through its respective encoder. For example, in VisualBERT[28], the first N embeddings represent textual information, while the following K embeddings represent visual information, separated by a [SEP] token. This allows the model to process and interpret the fused multimodal data jointly. Building upon this, LXMERT[29] enhances the model’s capacity to extract both intra-modal and inter-modal features by incorporating self-attention and cross-attention mechanisms. UNITER[30] further improves multimodal interaction learning by combining four training objectives: Masked Language Modeling (MLM), Masked Region Modeling (MRM), Word-Region Alignment (WRA), and Image-Text Matching (ITM).
Figure 7: Illustration of unified multimodal representation embedding methods.
In the field of remote sensing, RSGPT[31] integrates visual and language modalities using a Q-Former to fuse modality-specific embeddings. It introduces image-question-answer triplets based on remote sensing imagery interpretation as domain-specific knowledge, and proposes RSIEval as a benchmark for evaluating the performance of remote sensing multimodal large language models (RSMLLMs). Building on this, GeoChat[32] fine-tunes a pretrained LLaMA[16] model using four types of visual-language understanding tasks. It distinguishes visual and language modalities using <im_start> and <im_end> tokens, significantly improving the model’s understanding of remote sensing imagery. Similarly, RS-LLaVA[9] adopts the multimodal architecture of LLaVA[58], introducing a projection layer to map visual modalities. Remote sensing imagery and their descriptions are encoded as image tokens and text tokens, respectively, and the model is further optimized using LoRA[70] fine-tuning techniques.
Dynamic Hybrid Embedding
As shown in Fig.7(b), dynamic hybrid embedding refers to the approach in which data from different modalities are concatenated and directly input into the embedding layer for joint processing. Unlike the modality-tagging method, in dynamic hybrid embedding, the number and positions of embeddings from each modality are dynamically variable. This allows the model to effectively capture interaction patterns across modalities, thereby enhancing its understanding of multimodal joint representations. For example, OSCAR[33] directly combines textual data, object labels, and object images into a single input, which is uniformly processed through the embedding layer. This approach significantly improves the model’s ability to extract target information in query-based tasks. PaLM-E[35] builds on this by integrating continuous data streams from the physical world, enabling the model to interact with real-world environments. VIMA[34], focused on the domain of embodied intelligence, directly combines text tokens, object tokens, and action tokens, allowing the model to learn how to generate appropriate action feedback in response to different instructions and goals—thus enabling robots to perform a wide variety of tasks.
In the field of remote sensing, TEOChat[36] employs a Shared Vision-Language Connector to achieve hybrid embedding of image tokens and text tokens, further enhancing the fusion of multimodal information. To improve the model’s understanding and reasoning ability with respect to geographic knowledge, SkySenseGPT[37] incorporates more detailed instructions into the text tokens and introduces both region-level and image-level scene graph information. This significantly strengthens the model’s ability to understand both local and global relationships among geo-objects. Through this approach, remote sensing models are better equipped to capture spatial information and object relationships across different scales and levels, thereby improving their overall capability for intelligent interpretation of remote sensing imagery.
The shift from modality-tagging methods to joint embedding strategies has become a major development trend for improving the performance of language-centered multimodal models. The core idea is to increase the model’s learning difficulty by disrupting the order of tokens from different modalities, thereby encouraging deeper understanding of multimodal data. Modality-tagging methods assign fixed positions and lengths to different modalities, allowing the model to clearly distinguish the source of each token. However, this structured design can limit the model’s ability to capture complex relationships across modalities. In contrast, joint embedding mixes and concatenates tokens from different modalities in a less constrained manner, compelling the model to learn inter-modal interactions and complementary information within a more flexible and intricate input structure. As a result, the model must not only understand the individual characteristics of each modality but also learn how to integrate latent relationships across modalities—thus enhancing its overall capability to interpret multimodal data holistically.
III-B2 Semantic Alignment Across Modalities
Aligning representations from different modalities at the semantic level is a fundamental challenge in multimodal alignment research. Existing approaches to semantic alignment can be broadly categorized into two types: explicit alignment and implicit alignment. The key difference between them lies in whether the alignment relationships across modalities are explicitly introduced during training—such as through additional manual annotations or pre-defined alignment schemes.
Explicit Alignment
To effectively embed different modalities into the language modality, a straightforward approach is to use paired data—such as image-text or point cloud-text datasets—to train the model. Within this training framework, the model learns to align modalities by modeling the pairing relationships between multimodal data and corresponding textual descriptions. Based on the granularity of alignment between non-linguistic and linguistic modalities, current research efforts can be broadly classified into three main approaches: Scene-Caption alignment, Region-Phrase alignment, and Scene-Instruction alignment. The characteristics of each approach are illustrated in Fig.8.
Figure 8: Illustration of explicit alignment methods.
  • Scene-Caption Alignment: Scene-Caption alignment methods focus on aligning an entire image with its corresponding textual description, aiming to help the model learn the overall association between visual content and language, as illustrated in Fig.8(a). A representative work in this category is CLIP[7], which leverages contrastive learning to achieve effective alignment and has inspired a series of follow-up methods, including UniCL[38], K-LITE[39], and LaCLIP[40]. Due to the task-agnostic nature of contrastive learning, these approaches are well-suited as foundational models that can be quickly adapted to various downstream tasks via fine-tuning, few-shot, or even zero-shot learning. However, these methods face several limitations. First, they rely heavily on large-scale, high-quality image-text paired datasets, and their image understanding capability tends to be shallow. Second, such methods often struggle to handle both image understanding and text generation tasks simultaneously. To address these issues, SimVLM[41] introduced the PrefixLM pretraining objective, which maintains strong image understanding while enabling text generation capabilities. Similarly, BLIP[20] performs well in both image understanding and captioning, and its unique CapFilt mechanism effectively reduces noise in the training data. Despite these improvements, Scene-Caption-based methods generally lack the ability to deeply and finely understand image content. This limitation largely stems from the fact that simple image-text pairs fail to provide sufficiently rich guidance, thereby constraining the model’s capacity to learn more complex vision-language relationships.
    In remote sensing, vision-language models have rapidly become a research focus due to their superior capabilities in image understanding and language processing. One of the early efforts in Scene-Caption alignment for remote sensing was conducted by Qu et al.[42], who utilized the UCM Captions dataset. However, this dataset was relatively small in scale and contained simple descriptions, limiting its ability to represent complex content and diverse scenarios in remote sensing imagery. To address these limitations, subsequent works such as RSICD[43] and RSITMD[44] constructed larger, more diverse, and semantically richer datasets, while also developing task-specific models for various downstream applications. Building upon these foundations, RemoteCLIP[45] and GeoCLIP[46] have further advanced the field. These studies introduced large-scale Scene-Caption datasets for remote sensing containing up to 1.6 million and 5 million images, respectively, and proposed two foundational models tailored specifically for the remote sensing domain. Compared to earlier efforts, these studies adopt Transformer-based architectures, demonstrating a deeper understanding of remote sensing image content. As a result, they have established new directions and benchmarks for vision-language model research in remote sensing.
  • Region-Phrase Alignment: To enhance the model’s fine-grained understanding of images, some studies have explored incorporating detailed correspondences between image regions and individual words or phrases, as illustrated in Fig.8(b). Early work in this research direction—such as ViLBERT[47] typically relied on external object detectors to identify regions of interest (ROIs), using techniques like RoIAlign to extract features from these regions and align them with corresponding text tokens. However, this approach has several notable limitations. First, its dependence on pre-trained external detectors restricts the range of detectable object categories and directly impacts model performance. Second, the inclusion of an additional detection module increases model complexity and reduces training efficiency. To address these issues, VL-BERT[49] integrates the detection module into the training process, while ViLT[50] eliminates it altogether, instead using a pure Transformer architecture to extract image features. These adjustments significantly improve efficiency while maintaining strong performance. Additionally, some researchers argue that relying solely on ROI features may lead to the loss of valuable contextual information in the background. In response, strategies such as background blurring, masking, and feature fusion have been proposed to preserve as much background information as possible while extracting regional features. These strategies further enrich the model’s understanding of image content, especially when dealing with complex scenes containing rich background and multi-layered information.
    In remote sensing, Region-Phrase alignment has emerged as a compelling research direction, particularly within the context of Scene-Caption tasks. Early studies primarily adopted traditional encoder-decoder architectures, focusing on extracting features from target regions for image caption generation. While these approaches proved effective, they often overlooked domain-specific knowledge unique to remote sensing. With the rapid advancement of vision-language models, concept-level Region-Phrase alignment has seen new breakthroughs in remote sensing. For instance, GeoVG[55] was the first to introduce the Visual Grounding task into the remote sensing domain. It employs a dual-stream architecture to separately extract visual features and geographic relational graphs, which are then integrated through a fusion module to effectively combine geospatial relationships with visual information. Building upon this, RSVG[56] introduces the MGVLF module, which enhances grounding accuracy by filtering out noise and reinforcing salient features. Additionally, Few-Shot Object Detection (FSOD) has become an emerging topic in remote sensing. A recent work, TEMO[57], has made innovative contributions in this direction by incorporating detailed descriptions of all object classes as additional features, effectively mitigating class confusion. These advancements not only demonstrate the significant potential of Region-Phrase alignment in the remote sensing domain, but also provide valuable insights and directions for future research.
  • Scene-Instruction Alignment: While the previously discussed Scene-Caption and Region-Phrase alignment methods effectively achieve direct alignment between visual and language modalities, they often fall short in exploring the complex and abstract relationships among concepts expressed across modalities—an aspect that is crucial for achieving higher-level compositional generalization. Against this backdrop, Scene-Instruction-based alignment demonstrates its unique value. Its core lies in answering questions through logical reasoning, as illustrated in Fig.8(c). This requires the model not only to understand individual concepts but also to grasp the intricate relationships between them. In recent years, LLMs have demonstrated powerful general knowledge understanding and reasoning capabilities, which has driven a new research direction that leverages LLMs to tackle multimodal tasks. By fine-tuning LLMs with Scene-Instruction datasets—as done in LLaVA[58] and InstructBLIP[59] models can further enhance their logical reasoning and instruction-following abilities. These works have also simplified the process of constructing high-quality instruction datasets. Meanwhile, recent studies have introduced novel instruction processing mechanisms: for example, CAT[60] expands narrow language instructions into vision-language instructions, while SEAL[61] incorporates a visual search mechanism to help the model focus more precisely on key regions within an image. These innovations not only improve the model’s ability to understand complex scenes but also contribute new technical pathways toward the development of general artificial intelligence.
    In remote sensing, Scene-Instruction alignment is gradually emerging as a critical technique. It enables non-expert users to interact with remote sensing imagery using natural language. In RSVQA[62], researchers used OpenStreetMap data and manual annotation to create the first large-scale remote sensing visual question answering (VQA) dataset, and proposed a CNN-RNN-based VQA model. Subsequent work advanced the field by developing VQA-TextRS[63], a dataset supporting open-ended question answering, alongside a Transformer-based VQA model. In the same year, Prompt-RSVQA[64] proposed a method for mapping images into the language space and feeding them, along with questions, into BERT[71] to generate answers. These studies indicate that remote sensing research is rapidly converging toward the direction of vision-language models, laying the groundwork for future innovations in human-AI interaction and multimodal scene understanding.
In summary, the progression from Scene-Caption to Region-Phrase, and more recently to Scene-Instruction alignment reflects a gradual deepening in the semantic understanding of geo-objects within remote sensing intelligent interpretation. This evolution corresponds to a shift from coarse- to fine-grained image-text matching strategies. Specifically: Scene-Caption alignment focuses on matching entire images with holistic textual descriptions, representing a coarse-to-coarse granularity alignment. While this approach captures the overall semantics of an image, remote sensing imagery often contains multiple geo-objects, and global descriptions may struggle to adequately and precisely express the visual content present in the image. Region-Phrase alignment refines the image side by dividing it into multiple regions and aligning each with a corresponding phrase. Here, the image is fine-grained while the text remains relatively coarse (describing entire regions). This method allows for more accurate localization and description of specific geo-objects in the image, enabling multi-object interpretation. However, it lacks detailed semantic modeling of conceptual relationships between geo-objects. Scene-Instruction alignment further enhances the granularity of image-text alignment by using complex instructions to describe the relationships among geo-objects and guide the model’s attention across multiple objects in the image. This approach not only aligns multiple targets in the imagery but also encodes rich semantic instructions that express inter-object relationships, thereby deepening the model’s interpretive understanding of remote sensing scenes. The development of image-text alignment methods—from coarse to fine granularity—has accelerated the evolution of remote sensing interpretation models from shallow semantic understanding toward deep mechanistic comprehension.
Implicit Alignment
While explicit alignment methods have achieved remarkable success in multimodal learning, they also present several practical limitations. First, they rely on large-scale paired image-text datasets for model training, which poses significant challenges in terms of data collection and preprocessing. Second, achieving alignment between two different modalities typically requires separate training procedures for each modality, increasing both computational and time costs. More importantly, as the number of modalities increases, even when using a language-centered alignment strategy, the alignment process becomes exceedingly complex and cumbersome. Moreover, when interaction is required between two non-language modalities, using language as an intermediary can lead to issues in both efficiency and accuracy. If the alignment between one modality and language is imperfect, errors may accumulate during the conceptual transfer process, ultimately leading to suboptimal or incorrect interaction outcomes.
Despite the significant heterogeneity across different modalities, they ultimately describe the same physical world. This fundamental fact implies that, regardless of whether the alignment between visual and language modalities is explicitly defined, there should exist an intrinsic condition for unified convergence. In support of this, Minyoung Huh et al.[72] experimentally demonstrated that visual models with different architectures and training objectives exhibit a high degree of representational consistency, and that this consistency continues to improve as model size and training data scale increase. This pattern also holds for cross-modal vision-language models. Similarly, Maniparambil[65] showed that it is possible to perform subtitle matching or retrieval tasks using unaligned visual and language encoders—further validating the existence of an inherent consistency between visual and language representation spaces. These findings raise a fundamental question: Is there an implicit, inherently defined alignment between language and vision? If so, then it may be possible to achieve concept-level alignment between vision and language even when training on single-modal data—this is the central idea behind implicit alignment, which has recently garnered increasing research interest. Notably, many of these studies choose to align visual representations to the language space, rather than the other way around. This leads to a second key question: Why is language often chosen as the convergence point for vision-language alignment? This paper argues that, as the primary medium of human communication, language most accurately reflects human information cognition and logical reasoning mechanisms. Furthermore, language is inherently abstract, generalizable, and adaptable, making it well-suited as the global workspace for efficiently organizing a conceptual space that encompasses visual and other non-linguistic modalities. Building on this view, this paper refers to such research efforts as language-centered implicit multimodal alignment.
Language-centered implicit multimodal alignment aims to explore how non-linguistic modality information can be directly aligned with the language representation space in the absence of prior modality-language alignment pairs. In this line of research, the work LM4VisualEncoding[66] discovered that even when trained solely on textual data, the encoder of a large language model demonstrates surprisingly strong performance on both pure visual tasks and multimodal tasks. This finding offers new insights into the use of language models for visual tasks, highlighting their ability to encode and process visual-language information through shared parameters. Furthermore, the recent LAMP[67] study confirms the potential of pretrained language models in handling 3D point cloud data, providing additional evidence of the broad adaptability of language models in interpreting and processing non-traditional textual data. Collectively, these research efforts have pushed the boundaries of applying large language models to computer vision tasks and expanded our understanding of their latent capabilities in processing unimodal visual data.
Figure 9: Structural organization of language-centered knowledge association.
With the rapid advancement of remote sensing and related technologies, vast amounts of remote sensing data are generated daily. However, constructing large-scale image-text paired datasets in the remote sensing domain remains a significant challenge. Therefore, introducing implicit alignment into remote sensing is both necessary and timely. Recently, several studies have begun to explore this direction. For example, GRAFT[68] leverages internet images captured from the same ground locations as remote sensing images to serve as an intermediary between remote sensing imagery and language. It proposes a method to train a remote sensing vision-language model without relying on any textual annotations. Such attempts to introduce implicit alignment into the remote sensing field provide promising new directions for addressing the shortage of large-scale multimodal image-text pairs. Nonetheless, this line of research is still in its early stages and requires further exploration in the future.
III-C Language-Centered Knowledge Association
Alignment enables the convergence of information from different modalities into a shared language embedding space, whereas the core of knowledge association lies in establishing associations between perceptual information and internal/external knowledge based on the interpretation task requirements. Accordingly, the knowledge association process primarily involves two key challenges:
  • How to understand the interpretation task?
  • How to identify the relevant knowledge?
The structure of this section is illustrated in Fig.9, and representative methods for each component are summarized in Table II.
TABLE II: Language-Centered knowledge association structure and representative methods.
Content structure Category Representative algorithms
Task Understanding Rule-Based APLICOT[73], ACT-R[74], RS-DFM[75], Feng et al.[76], Semantic-CD[77]
Pretrained Language Models Based Llama3[78], GPT-3[17], DeepSeek-V3[79], LiT-4-RSVQA[80], GeoGPT[81], GeoChat[32], SkyEyeGPT[82], EarthGPT[10]
Reasoning-Based CoT[26], Self-Reflection[83], Self-Refine[84], OpenAI-O1[85], DeepSeek-R1[27], Remote Sensing ChatGPT[86], CEOS[87]
Knowledge Base Driven Auto-CoT[88], MemoryBank[89]
Physical World Interaction Driven ReAct[90], ReSpAct[91], RS-Agent[92]
Knowledge Retrieval Static Knowledge Encoding Word2Vec[93], GloVe[94], FastText[95]
Dynamic Knowledge Encoding ELMo[96], BERT[71], GPT[97]
Multimodal Joint Encoding MKBE[98], MULE[99], E5-V[100], EarthMarker[11]
Static Vector Search TF-IDF[101], k-NN[102], ANN[103]
Dynamic Contextual Reasoning Self-attention[104], Cross-attention[105], Block-Attention[106], KTIR[107]
III-C1 Task Understanding
The depth of task understanding in LLMs is reflected in their ability to decompose complex tasks. For example, a ship motion monitoring task requires the interpretation system to detect and recognize all ships in a specified maritime region, identify their fine-grained categories, and determine dynamic changes over time through temporal analysis. Due to the highly composite nature of this task, a single model would struggle to handle it effectively. Instead, an LLM is capable of decomposing it into four relatively independent yet sequentially connected subtasks: detection, coarse classification, fine-grained recognition, and temporal analysis each of which can be delegated to a specialized expert model. In this process, the LLM effectively re-structures the complex task of ”ship motion monitoring” into a series of simpler subtasks, each with a clear knowledge objective and a corresponding expert model. This demonstrates the LLM’s comprehension of the task and also reduces the overall task complexity, thereby facilitating more efficient task execution. From this perspective, based on whether the task environment including the knowledge space and the external environment changes during task understanding, we categorize task understanding into two types: Closed-Form task understanding, and Open-Ended task understanding.
Closed-Form Task Understanding
In closed-form task understanding scenarios, both the knowledge space and the environment space are static and fixed. Task understanding is primarily achieved through existing knowledge, without involving interaction with the external environment. There are three main approaches to task decomposition in this setting: Rule-based task decomposition, Pretrained language model-based task decomposition, and Chain-of-thought-based task understanding.
Figure 10: Illustration of closed-form task understanding methods.
  • Rule-Based Task Decomposition: Rule-based task decomposition relies on predefined rules and fixed logical procedures to break down and execute tasks, as illustrated in Fig.10(a). These rules are typically handcrafted for specific task requirements. Once the model receives a task input, it follows the established rules step by step to reason and complete the task. This approach offers strong interpretability, as each reasoning step and decision-making rule is explicit and fixed, allowing the process by which the model derives its output from the input to be clearly traced, understood, and verified. In the field of natural language processing, early rule-based systems such as APLICOT[73] (based on Prolog) used logical rules (e.g., A :- B, C, meaning ”IF B AND C, THEN A”) to connect multiple rule units and perform reasoning in complex tasks. However, this approach suffered from limited context understanding, as it could only build logical relationships based on local information. To address this shortcoming, ACT-R[74] extended APLICOT[73] by introducing a cognitive framework that combines fixed logic rules with memory units, enabling higher-level cognitive processing and improving flexibility and efficiency in task decomposition. Although rule-based methods can effectively support task understanding and decomposition under controllable logical chains, their reliance on explicitly defined rules has become a significant limitation in the era of big data. When dealing with complex tasks, the rigidity of rules often leads to poor performance. As a result, with the rise of language model-based methods, rule-based decomposition has been gradually phased out particularly due to language models’ superior flexibility and adaptability in handling diverse task scenarios.
    In the field of remote sensing, complex tasks often rely on expert prior knowledge, with predefined rules used to decompose tasks into subcomponents, which are then processed step by step in conjunction with interpretation models. For example, in the area of target recognition, RS-DFM[75] addresses 3D object detection under multi-platform observation conditions by decomposing the task into four fixed subtasks: feature extraction, BEV (bird’s-eye view) generation, high- and low-frequency feature decoupling and calibration, and feature fusion, in order to improve detection accuracy and robustness. In multimodal interpretation, Feng et al.[76] (2023) break down cross-modal interpretation into two subtasks: multi-scale spatial feature extraction and multimodal feature fusion. In change detection, Semantic-CD[77] decomposes the task into semantic segmentation and change feature extraction. It uses an Adapted Vision Encoder to extract semantic features from bi-temporal images, followed by a semantic change detection decoder to identify areas of semantic variation. These methods demonstrate that rule-based task decomposition plays an important role in remote sensing interpretation, as it not only simplifies complex tasks but also enhances accuracy and operational stability. However, such approaches are heavily dependent on expert-designed priors, with fixed task decomposition structures, making them poorly suited for open and dynamic environments. Their inability to adapt subtask structures according to changing scenes limits their flexibility and generalizability in complex task settings.
  • Pretrained Language Models Based Task Decomposition: As illustrated in Fig.10(b), task decomposition based on pretrained language models leverages large-scale models such as Llama3[78], GPT-3[17], and DeepSeek-V3[79]. These models are trained on massive text corpora and possess the ability to understand and flexibly handle a wide range of tasks. Task decomposition in this paradigm is typically performed by interpreting task descriptions or instructions provided as input, and generating output based on the model’s pretrained knowledge and contextual understanding. The primary advantage of this approach lies in its flexibility: pretrained language models can handle a wide variety of tasks and adapt to diverse input types without the need for manually designed rules or constraints. Thanks to their strong language understanding capabilities, these models can provide reasonable solutions across different scenarios and show superior adaptability to changing task requirements. However, this flexibility comes at a cost—particularly in terms of interpretability. Since the decision-making process of these models is based on internal parameters and massive training data, the reasoning steps are often opaque and difficult to verify externally. As a result, the outputs may be non-traceable and lack guaranteed correctness.
    In the remote sensing domain, as new tasks continue to emerge—such as Remote Sensing Visual Question Answering (RSVQA), Remote Sensing Visual Captioning (RSVC), and Remote Sensing Visual Grounding (RSVG)—interpretation tasks are increasingly expressed in natural language. Rule-based task decomposition methods struggle to adapt to this flexible task format, prompting researchers to explore how to more efficiently ”understand” remote sensing interpretation tasks. To this end, several studies have introduced LLMs to analyze task requirements. For example, LiT-4-RSVQA[80] uses the text encoder of a VLM to interpret task instructions, while GeoGPT[81] employs an LLM to analyze and decompose interpretation tasks. Moreover, works such as GeoChat[32], SkyEyeGPT[82], and EarthGPT[10] further integrate LLMs into interpretation systems to enhance their understanding and decomposition capabilities for complex tasks. While LLM-based task decomposition offers significantly greater flexibility, allowing the system to automatically adapt to different task demands and intelligently break down complex tasks, the lack of transparency in the reasoning process introduces uncertainty regarding the trustworthiness of the decomposition results. This issue limits the practical applicability of such approaches in remote sensing scenarios that require high reliability.
  • Reasoning-Based Task Decomposition: Reasoning-based task decomposition methods combine the flexibility of language models with the interpretability of rule-based systems by guiding the model to perform step-by-step logical reasoning, as illustrated in Fig.10(c). In this approach, reasoning instructions are provided to the model in the form of prompts, alongside the input query. During the generation of subtasks, the model continuously rethinks and refines its outputs, effectively decomposing the task through a progressive reasoning process. This method encourages the model to self-check at each step, ensuring that task execution is both structured and logically sound. The earliest representative of this approach is the Chain-of-Thought (CoT)[26] prompting method, which uses simple cues like ”Let’s think step by step” to guide the model in generating intermediate reasoning steps and clearly showcasing the decomposition process in the final result. However, the reasoning logic in CoT[26] can be difficult to control, leading to unstable outcomes. To address this, methods such as Self-Reflection[83] and Self-Refine[84] introduce reflection mechanisms, enabling the model to self-evaluate the logic and accuracy of its reasoning, thereby improving reliability. Still, these methods rely on manually crafted prompt templates, and their performance on more complex tasks remains limited. To overcome these challenges, OpenAI-O1[85] and DeepSeek-R1[27] apply reinforcement learning to fine-tune the reasoning capabilities of language models. By integrating self-verification into the training process, these models are able to automatically analyze task steps without relying on external prompts, significantly enhancing their ability to understand and solve complex tasks with greater stability and precision.
    In the remote sensing domain, similar approaches have been explored in Remote Sensing ChatGPT[86], which interprets user instructions in natural language and combines them with domain-specific geoscientific knowledge to decompose complex interpretation tasks into subtasks. These subtasks are then handled sequentially by invoking corresponding expert models. This strategy not only enhances the automation of remote sensing interpretation, but also reduces dependence on domain experts. Furthermore, CEOS[87], developed based on the LangChain[108] framework, serves as an Earth observation analysis platform capable of automatically executing tasks such as vegetation analysis. These works demonstrate the potential of combining logical reasoning and task decomposition to improve the efficiency of remote sensing interpretation. However, these methods often rely on general-purpose LLMs guided by CoT-style prompting, and still suffer from inaccuracies when reasoning over specialized geoscientific knowledge. Due to the complexity of remote sensing imagery and the high level of expertise required in geoscience, existing general LLMs lack sufficient domain-specific knowledge and precise reasoning abilities. Currently, efforts to enhance LLMs’ geoscience reasoning through reinforcement learning-based fine-tuning are still in their early stages.
In summary, while closed-task understanding methods offer certain advantages in task decomposition, they also face inherent limitations. Due to the static nature of the knowledge and environment spaces involved in these methods, models are heavily reliant on predefined rules or pretrained knowledge—which often fall short when confronted with dynamic environments or unseen task requirements. This limitation is particularly evident when dealing with tasks that fall outside the coverage of the model’s internal knowledge base, where LLMs often struggle to deliver satisfactory results. Although rule-based and pretrained language model-based methods can effectively handle well-defined tasks and achieve a high level of automation, their adaptability and flexibility are constrained by the static nature of their knowledge and the lack of interpretability in their reasoning processes. Furthermore, while reasoning-based task decomposition methods attempt to enhance reasoning robustness through chain-of-thought prompting and reflection mechanisms, they still encounter difficulties when facing knowledge gaps or complex task scenarios. As a result, closed-task understanding methods continue to face significant challenges in terms of performance and reliability when applied to dynamic and complex real-world tasks.
Open-Ended Task Understanding
In open-ended task understanding scenarios, both the knowledge space and the environment space can be dynamically adjusted according to actual task requirements. To overcome the inherent limitations of closed-task understanding, researchers have begun to explore the integration of external information into the task understanding process. Based on the source of external information, open-task understanding can be further divided into two main pathways: Knowledge base–driven task understanding, and Physical world interaction–driven task understanding.
Figure 11: Illustration of Open-Ended Task Understanding Methods.
  • Knowledge Base–Driven Task Decomposition: As illustrated in Fig. 11(a), knowledge base–driven open-task understanding enhances a language model’s comprehension by enabling interaction with external knowledge bases. The model retrieves information from these sources that is similar or relevant to the input query, thereby injecting additional knowledge into the task decomposition process. Representative works include Auto-CoT[88], which retrieves task decomposition examples similar to the input instruction and guides the LLM to imitate these examples for more reasonable task breakdowns. Another example is MemoryBank[89], which leverages Retrieval-Augmented Generation (RAG) to extract task-relevant knowledge from an external database and supports the decomposition process accordingly. Because external knowledge bases can be updated in real time, this method allows LLMs to continuously access the latest information and fill knowledge gaps as needed. However, the relevance and usefulness of the retrieved knowledge still depend heavily on the accuracy of the retrieval algorithm, which determines whether the acquired information truly aligns with the task requirements.
  • Physical World Interaction–Driven Task Decomposition: As illustrated in Fig.11(b), physical world interaction–driven open-task understanding acquires the required information through feedback from interactions with the physical environment, enabling the task decomposition process to unfold progressively. The language model initially generates a set of subtasks based on the input query; as actions associated with these subtasks affect the physical world, feedback from the environment is used to refine and optimize subsequent decomposition steps. A representative work in this direction is ReAct[90], which combines task decomposition with action execution. By directly interacting with the physical world, the model determines the next steps of task decomposition based on both environmental feedback and the results of previous task planning. ReSpAct[91] extends this framework by introducing user interaction feedback, allowing the LLM to seek advice, clarify ambiguities, or understand user preferences through dialogue—thereby updating its task decomposition plan in a more responsive manner. This approach—where task understanding is achieved through direct interaction with the physical environment—enables LLMs to dynamically adjust decomposition strategies according to changes in the environment, thus providing greater adaptability and flexibility. However, a key challenge in this line of research remains: how to design efficient and reliable environment interaction mechanisms.
In the field of remote sensing, open-task understanding is still in its early stages of development. One of the most recent relevant studies is RS-Agent[92]. RS-Agent[92] builds upon a pretrained LLM and implements five specialized agents responsible for reasoning, decision-making, memory, knowledge-enhanced retrieval, and decision-enhanced retrieval, respectively. It also establishes integration channels with conventional remote sensing interpretation tools such as object detection models, superpixel reconstruction models, scene classification models, and change detection models.
Figure 12: Illustration of knowledge encoding methods.
This architecture enables the model to retrieve task-relevant knowledge from a broader external knowledge base, thereby improving its capacity for task understanding. However, RS-Agent[92] relies solely on external knowledge bases for information retrieval and lacks mechanisms for interacting with and receiving feedback from the real physical environment. As a result, the model struggles to assess whether task decomposition is accurate, appropriate, and efficient under dynamic and open-world conditions. Thus, enabling effective interaction and feedback with the physical environment will be a critical challenge for the future development of RS-Agent[92] and similar research efforts.
III-C2 Knowledge Retrieval
Once the task requirements have been parsed and clarified during the task understanding process, the core challenge shifts to how to extract highly relevant information from internal and external knowledge bases to support subsequent reasoning and decision-making. Two fundamental components of knowledge retrieval are knowledge encoding and knowledge querying. Knowledge encoding focuses on transforming external knowledge into representations that the model can understand and utilize effectively in later stages. Knowledge retrieval, on the other hand, is guided by the specific content of the task requirements and aims to selectively extract information that is highly relevant from both internal and external sources. This ensures the model has access to accurate and timely knowledge, which is critical for effective reasoning and task execution. Through precise knowledge retrieval, the model can compensate for its internal knowledge gaps and leverage existing knowledge bases to further improve its capability and accuracy in handling complex tasks.
Knowledge Encoding
External knowledge is typically stored in the form of data, and the process of knowledge encoding essentially involves transforming this multimodal data into vector representations and mapping it into a unified vector space to facilitate subsequent knowledge association and retrieval. Within this space, semantically related knowledge should be located close together, while irrelevant knowledge should be distant. The core objective of this process is to enable the model to efficiently search, access, and integrate relevant knowledge. With the advancement of knowledge retrieval technologies, knowledge encoding has evolved through three key stages: Static encoding, Dynamic encoding, and Multimodal joint encoding.
  • Static Knowledge Encoding: As illustrated in Fig.12(a), static knowledge encoding maps different words into a shared semantic space where semantic relationships between words are preserved—for example, between ”Woman” and ”Queen”, or ”Queen” and ”King”. One of the pioneering methods in this domain is Word2Vec[93], which effectively captures both semantic and syntactic relationships between words. It uses two architectures CBOW (Continuous Bag of Words) and Skip-gram to project high-dimensional word representations into a lower-dimensional vector space, ensuring that semantically similar words are positioned closer together. However, a major limitation of Word2Vec[93] is its inability to capture the dynamic semantic shifts of words across different contexts. To address this, GloVe[94] proposes a matrix factorization approach based on word co-occurrence statistics, allowing it to encode global semantic relationships more effectively. Building on this, FastText[95] refines word representation further by breaking words down into subword units. This enables the model to learn internal word structures, thereby improving its ability to represent low-frequency or out-of-vocabulary words. Nevertheless, static knowledge encoding fails to reflect contextual semantic variation. As shown in Fig.12(a), methods like Word2Vec[93] tend to map the word ”bank”—regardless of whether it refers to a financial institution or a riverbank—to the same point in the semantic space, leading to semantic ambiguity.
  • Dynamic Knowledge Encoding: Unlike static encoding methods, which assign a fixed vector representation to each word, dynamic knowledge encoding generates context-dependent word representations, enabling more accurate semantic interpretation, especially in the presence of contextual variability and polysemy. As illustrated in Fig.12(b), the word ”bank” can be mapped to different semantic regions depending on its surrounding context—whether referring to a financial institution or a riverbank. A representative method in this category is ELMo[96], which utilizes a pretrained bidirectional language model to produce dynamic word embeddings based on surrounding context, effectively capturing the context-specific meanings of words. BERT[71] takes this a step further by employing the bidirectional self-attention mechanism of the Transformer architecture, allowing it to simultaneously consider both the left and right context of a word to generate highly context-aware embeddings. In contrast to BERT[71], GPT[97] is a unidirectional language model trained to predict the next token in a sequence. This training strategy enables GPT[97] to compress semantic information from the input corpus efficiently, despite not incorporating full bidirectional context.
  • Multimodal Joint Encoding: While dynamic knowledge encoding has made significant progress in handling language-based knowledge representation, it struggles to effectively process multimodal data. In contrast, multimodal knowledge encoding offers a more comprehensive and unified solution for integrating information across diverse modalities. As illustrated in Fig.12(c), this approach enables the joint representation of vision, language, and other modalities to describe a shared semantic concept, laying a solid foundation for downstream knowledge retrieval and reasoning. In this field, MKBE[98] constructs a multimodal knowledge base by unifying the representations of visual and language modalities. During the construction process, MKBE[98] focuses on preserving temporal characteristics of each modality and the relationships between entities expressed across modalities. For example: For the visual modality, MKBE[98] uses Convolutional Neural Networks (CNNs) to extract spatial and semantic features from images. For the language modality, LSTM[109] networks are employed to capture sequential dependencies between words. At the entity-relation level, MKBE[98] explicitly expresses cross-modal relationships by designing subject-relation pairs as training ground truth, and uses a cross-entropy loss function to train the model to recognize entity relationships. Building on MKBE[98], MULE[99] further enriches the language modality by supporting multilingual representations, thereby improving the model’s compatibility with multiple languages. Although these methods have made notable advances in multimodal joint encoding, they still require dedicated embedding modules for each individual modality. With the rapid development of LLMs, newer approaches like E5-V[100] leverage MLLMs as unified encoders, allowing different modalities to be directly mapped into a shared embedding space without the need for modality-specific components. This approach significantly simplifies the implementation of multimodal joint encoding and improves the efficiency of cross-modal data integration.
    In the field of remote sensing, the unified representation and understanding of multimodal data remains a central challenge for intelligent interpretation systems. The main difficulty lies in effectively fusing information from different modalities. With the rapid advancement of LLMs, a language-centered paradigm for multimodal joint embedding has gained increasing attention. EarthMarker[11] applies this paradigm to remote sensing by embedding visual modality data into the representational space of language models, thereby constructing a modality-unified semantic space. Within this framework, EarthMarker[11] introduces the MoV (Mixture of Visual Experts) and a modality alignment module to integrate remote sensing imagery, visual object prompts, and textual instructions. This enables unified understanding and representation of heterogeneous multimodal information. This approach not only enhances the interpretability of remote sensing data, but also offers a novel pathway for intelligent processing in multimodal tasks.
Figure 13: Illustration of knowledge retrieval methods.
In summary, knowledge encoding transforms external knowledge into vector representations that can be understood by models, laying the foundation for subsequent knowledge retrieval. The development of knowledge encoding has gone through three major stages—static encoding, dynamic encoding, and multimodal joint encoding—each enhancing the expressive power of knowledge in different ways. Static encoding methods, such as Word2Vec[93], GloVe[94], and FastText[95], focus on capturing semantic relationships between words. However, they are limited by their inability to dynamically adapt to changes in context. Dynamic encoding methods, such as ELMo[98], BERT[71], and GPT[97], can adjust word embeddings based on contextual information, allowing for more flexible and accurate semantic interpretation. Multimodal joint encoding goes a step further by integrating information from multiple modalities—such as vision and language—thus providing robust support for cross-modal knowledge retrieval and reasoning.
Knowledge Retrieval
Knowledge retrieval is a crucial step following knowledge encoding, aiming to extract information that is highly relevant to the current task from large-scale knowledge bases or databases. This information is then used to support subsequent reasoning and decision-making processes. With the advancement of technology, knowledge retrieval has evolved from traditional keyword-based search methods to more intelligent approaches, such as static vector-based retrieval and dynamic context-aware reasoning. These modern methods enable models to identify the most valuable and relevant knowledge from massive information repositories, even in complex task scenarios, thereby enhancing task comprehension and decision-making performance.
  • Static Vector Search: Static vector–based knowledge retrieval relies on precomputed knowledge embeddings. As illustrated in Fig.13(a), the system retrieves relevant knowledge by calculating the similarity between the query vector and stored knowledge vectors. Common techniques include TF-IDF[101], k-Nearest Neighbors (k-NN)[102], Approximate Nearest Neighbor (ANN)[103] search, and cosine similarity. TF-IDF[101] evaluates the importance of keywords by measuring their frequency in a document relative to their prevalence across the entire corpus. However, it ignores semantic relationships between words. The k-NN[102] algorithm finds the most similar knowledge vectors by computing distances between the query and the stored vectors. To improve efficiency, ANN[103] uses optimized indexing structures to perform approximate retrieval, reducing computational cost while maintaining acceptable accuracy. Cosine similarity measures the angle between query and document vectors, offering a scale-invariant similarity measure that avoids the influence of vector length. These methods are effective for static vector retrieval but lack adaptability and often struggle with complex semantic understanding tasks, where deeper context and reasoning are required.
    Figure 14: Structural organization of language-centered reasoning and decision-making.
  • Dynamic Contextual Reasoning: Unlike static vector search, dynamic context–aware knowledge retrieval (as illustrated in Fig.13(b)) leverages attention mechanisms[69, 105] to establish relationships between pieces of knowledge within a broader context, significantly improving retrieval precision in complex semantic understanding tasks. The attention mechanism operates by computing attention scores through interactions among queries, keys, and values, enabling the model to identify the most relevant knowledge across the global context. In traditional attention frameworks, the query is matched against all key-value pairs in the knowledge base. However, as the number of knowledge vectors grows, the computational cost of this matching process increases dramatically, leading to long retrieval times. To address this issue, Block-Attention[106] introduces a block-wise strategy, dividing large query sequences into smaller chunks and performing attention calculations within each block independently. This approach significantly reduces computational overhead. For example, with a prompt length of 32K, Block-Attention reduces TTFT-block latency by 98.7%. By optimizing the attention-based retrieval process, this method improves the efficiency of dynamic knowledge retrieval in large-scale knowledge bases.
    In the remote sensing domain, effective text-to-image retrieval is a key challenge for Earth observation tasks. Due to the vast volume of remote sensing imagery and the minimal variance among image embeddings, traditional text retrieval techniques often struggle to match images accurately. To address this, the KTIR[107] (Knowledge-to-Image Retrieval) framework proposes a cross-modal retrieval method. It attributes poor retrieval performance to the lack of differentiation in textual descriptions. KTIR[107] addresses this by converting structured knowledge within remote sensing images into textual expressions, and then applying a Cross-Attention[105] mechanism that uses these expressions as queries for retrieval. This approach significantly improves the retrieval accuracy between remote sensing images and textual descriptions.
In summary, static vector–based retrieval methods perform information retrieval by calculating similarity scores between query and knowledge vectors, but they face significant limitations in capturing contextual variations. To overcome this challenge, dynamic context–aware knowledge retrieval methods leverage attention mechanisms to establish relationships across a broader context, thereby enhancing the accuracy of knowledge retrieval in complex semantic scenarios.
TABLE III: Corresponding representative methods of language-centered reasoning and decision-making.
Content structure Category Representative algorithms
Task Progress Assessment Behavior Cloning Thought Cloning[110], EBC-LLMAgent[111]
Inverse Reinforcement Learning MA-AIRL[112], Xirl[113]
Reward-Based Liir[114], ExpRD[115]
Policy Learning–Based PPO[116], DQN[117]
Model Predictive Control MPC[118], DMPC[119], L1-NMPC[120]
Decision Making Action-Level Decision Making Hu and Pagilla[121], Namiko[122], PID[123], LQR[124], DeepMPC[125], DeepGait[126], GDQ[127], RobotPilot[128], PIBOT[129]
Tool-Level Decision Making ReACT[90], DFSDT[130], Tool-bench[130], ToolLLaMA[16], Levine et al.[131], LangChain[108], AutoGPT[132], CLIPort[133], Toolformer[134], Re-Invoke[135], WebGPT[136], WebCPM[137]
Figure 15: Illustration of retrospective evaluation methods.
III-D Language-Centered Reasoning and Decision-Making
A core challenge in reasoning and decision-making is how to derive subsequent sensing strategies in a task-driven manner by combining perceptual information, associated knowledge, and available actions. This process involves inferring what to do next based on the current understanding of the task. There are two key problems in this reasoning and decision-making phase: How to assess the gap between external information and task objectives, and How to make decisions based on this gap. The structure of this section is illustrated in Fig.14, and representative methods are summarized in Table III.
III-D1 Task Progress Assessment
Assessing the gap between external information and task objectives is critical for making correct action decisions and satisfying interpretation requirements. In addressing this issue, some studies evaluate the gap by comparing the model’s behavior to expert demonstrations, while others assess it by predicting the potential outcomes of the model’s actions. Based on whether expert demonstrations are used, existing studies can be categorized into two main paradigms: Retrospective Assessment, and Predictive Assessment.
Figure 16: Illustration of predictive evaluation methods.
Retrospective Assessment
Retrospective assessment assumes that the gap between the current state and the task objective can only be determined after task execution is completed. This type of evaluation primarily relies on comparing the model’s behavior with expert demonstrations to assess task discrepancies. There are two main approaches: Behavior Cloning, and Inverse Reinforcement Learning (IRL). Both methods analyze the differences between the model’s behavior and expert behavior, helping to identify deficiencies in the model’s task execution and providing a foundation for further optimization.
  • Behavior Cloning: These methods assess the gap between task objectives and perceptual information by imitating expert demonstration behaviors, as shown in Fig.15(a). The key idea is to treat expert behavior as the learning target, allowing the model to analyze task progress by cloning expert actions. For example, Thought Cloning[110] mimics the human reasoning process, which is not only focusing on the actions themselves, but also modeling the decision-making logic behind them. It quantifies the discrepancy between the model’s behavior and expert goals at both the logical reasoning and execution levels. Similarly, EBC-LLMAgent[111] leverages interpretable mechanisms to link perceptual inputs (e.g., raw data) with expert decision intentions, enabling the model to evaluate how far its behavior deviates from the intended goals. These approaches provide a direct and interpretable pathway for assessing the gap between perceptual information and task objectives by explicitly constructing mappings between behavior and intent.
  • Inverse Reinforcement Learning: Unlike behavior cloning, this class of methods does not directly imitate expert behavior. Instead, it infers the underlying reward function behind expert actions to derive an optimal policy, as shown in Fig.15(b). This approach is particularly suitable for scenarios where expert demonstrations are limited or where designing explicit reward functions is challenging. In vision-language tasks, Inverse Reinforcement Learning (IRL) can be used to infer human decision-making preferences, enabling more reasonable gap assessments and improved model behavior. For example: MA-AIRL[112] employs Generative Adversarial Networks to model reward functions in multi-agent interactive environments, evaluating each agent’s contribution to global goals and quantifying the gap between perceptual information and task objectives. Xirl[113] infers reward functions across different environments to construct a generalizable reward representation for gap assessment. Its core idea is to use visual feature embeddings to achieve cross-environment alignment of perceptual information and use the learned reward function to reflect the deviation between perception and target behavior. Together, these two approaches offer distinct theoretical and practical frameworks for gap evaluation and policy optimization, effectively addressing the challenges posed by complex tasks.
Predictive Assessment
While retrospective assessment methods are effective for evaluating task gaps, they often rely on expert demonstrations, which can be difficult to obtain. To address this limitation, predictive assessment methods have emerged. These approaches aim to evaluate the gap between the current state and task objectives by predicting the model’s own future behavior, thereby determining whether upcoming actions are likely to lead to task success. Within this framework, predictive assessment methods can be further categorized into three types: Reward-based approaches, Policy learning–based approaches, and Model predictive control (MPC)–based approaches. By performing pre-execution predictions, these methods proactively identify potential execution issues and provide guidance to the model, helping it improve performance in real-world tasks.
  • Reward-Based: These methods design reward functions to quantify the gap between external information and task objectives, as shown in Fig.16(a), and use reinforcement learning algorithms to optimize policies that minimize this gap. For example: Liir[114] learns intrinsic rewards for individuals, combining their internal motivations with global objectives. In vision-language tasks, this mechanism enables personalized rewards for sub-goals, allowing for more fine-grained task gap evaluation. ExpRD[115] constructs interpretable reward functions that align reward design with human intuition, helping assess the discrepancy between model behavior and task goals. In such methods, the design of the reward function directly affects both the convergence speed and final performance of the model. Therefore, ensuring the rationality and flexibility of the reward function remains a critical research focus.
  • Policy Learning-Based: These methods optimize decision-making policies directly using policy gradients or value function approximation techniques, as shown in Fig.16(b). The goal is to maximize cumulative rewards, while evaluating task performance by quantifying the gap between the current policy’s return and the optimal policy’s return. Specifically: Value-based algorithms (e.g., DQN[117]) estimate Q-values to assess the potential deviation between current actions and target objectives. Policy-based algorithms (e.g., PPO[116]) optimize the policy distribution directly to narrow the reward gap. A core challenge in this approach is accurately modeling the relationship between the expected return of the current policy and the optimal return. Additionally, balancing exploration and exploitation is essential for improving learning efficiency and policy refinement in complex tasks.
  • Model Predictive Control (MPC): MPC–based methods evaluate the gap between task objectives and perceptual information by dynamically predicting future states, as illustrated in Fig.16(c). These methods emphasize real-time planning during the decision-making process, making them well-suited for scenarios that require continuous control and dynamic adjustment. By constructing a dynamic model, MPC[118] predicts the potential impact of current actions on future states and quantifies the deviation between actions and goals by optimizing a cost function (e.g., reconstruction error or goal matching score), thereby selecting the optimal action. For example: DMPC[119] uses a distributed control framework to decompose a complex global planning task into local optimization problems among multiple agents. It evaluates trade-offs between obstacle avoidance and task objectives, achieving coordinated global planning. L1-NMPC[120] applies adaptive nonlinear control to dynamically optimize the cost function, accurately quantifying the deviation between current actions and target states. It is particularly effective in high-precision control scenarios for gap evaluation and strategy optimization.While these methods demonstrate strong planning capabilities in complex and high-dimensional tasks, they typically involve high computational costs and rely heavily on the accuracy of the dynamic model.
In summary, retrospective assessment relies on expert demonstrations after task completion. It evaluates the model by comparing its behavior with that of experts through behavior cloning or inverse reinforcement learning, helping to identify model deficiencies and offering directions for optimization. However, this approach is limited by the high cost of acquiring expert demonstrations, making it difficult to scale in large-scale, dynamic tasks. In contrast, predictive assessment methods do not require expert demonstrations. Instead, they evaluate the potential impact of model behavior before execution, allowing for early estimation of the likelihood of task success. These methods can be further divided into reward-based, policy learning–based, and model predictive control–based approaches, offering a foundation for real-time strategy adjustments during task execution. Overall, retrospective assessment provides precise behavioral comparisons but depends heavily on expert data, while predictive assessment enhances task adaptability through self-prediction, though its accuracy and stability in prediction remain key challenges to be addressed.
III-D2 Decision Making
Once the gap between the current interpretation process and the task objective is clearly identified, the core challenge shifts to determining how to make informed decisions on the next interpretation actions, using task-relevant knowledge and the agent’s capabilities, in order to narrow that gap. This gap is primarily caused by insufficient perceptual information, which arises from two dimensions: the data level, and the functional level. To address these issues, it is necessary to adopt action-level decision-making methods and tool-level decision-making strategies, respectively.
Action-Level Decision Making
At the data level, the gap often manifests as the absence of key visual information required for interpretation in the collected imagery. This may result from limitations in observation conditions, image quality issues, or critical information not being captured during data acquisition. Action-level decision making addresses this issue by adjusting or controlling the sensor’s behavior based on perceptual feedback to supplement or improve the data collection process. This includes modifying sensor configurations, changing the observer’s position or orientation, or switching sensor modalities. For example: Hu and Pagilla[121] used perceptual feedback to control a robotic arm’s viewing direction and angle, enabling it to collect denser and higher-quality point cloud data for accurate pose estimation. Namiko et al.[122] combined multimodal sensor feedback with the robot’s motion status to make action decisions that accurately associate tools, objects, and actions.
In the remote sensing domain, action-level decision making is primarily applied to adjust UAV flight paths, altitudes, speeds, and viewing angles dynamically, adapting to changing environmental conditions and task requirements. Traditional action decision methods are based on classical control theory and dynamic modeling, typically using PID[123] controllers and LQR[124] controllers to stabilize UAV behavior and adjust observational states. While these approaches perform well in stability control, they often struggle in nonlinear, multivariable, and coupled dynamic environments. With the advancement of deep learning, Deep Reinforcement Learning (DRL) has become a prominent research direction in action decision-making. DeepMPC[125] combines reinforcement learning and MPC[118] to enable real-time pose adjustment in complex environments. DeepGait[126] integrates visual input and global positioning to allow quadruped robots to optimize their movement according to terrain conditions. GDQ[127] enhances multi-agent pose control in mapless navigation tasks by combining a global value network with Double Deep Q-Network (DDQN). In addition, imitation learning has shown great promise in this field. For example, RobotPilot[128] imitates human-operated UAV flight trajectories, achieving high-precision, task-oriented pose control. Recent research also highlights the potential of combining Vision-Language Models (VLMs) with pose estimation to improve UAV control. For instance, PIBOT[129] integrates a VLM with pose estimation capabilities to enable autonomous flight control and human-like pose adjustments for robotic pilots.
Tool-Level Decision Making
At the functional level, the gap stems from the mismatch between the agent’s decision outputs and the available remote sensing interpretation tools. The agent’s decisions are expressed in natural language, while interpretation algorithms and models exist as code modules. Establishing an effective mapping between language and code is the core challenge in aligning decision outputs with available tools. Tool-level decision making addresses this problem. LLMs can directly select and invoke tools by aligning tool descriptions with subtask descriptions expressed in natural language. ”Tuning-free” approaches, such as ReACT[90] and DFSDT[130], achieve this alignment through prompt engineering, guiding the LLM to map subtask instructions to tool descriptions. However, the ToolBench[130] study reveals that such methods often suffer from low tool selection accuracy. To overcome this, it proposes fine-tuning LLMs with tool descriptions. Following this idea, ToolLLaMA[16] re-trained the LLaMA-7B[16] model using an instruction-solution format to improve tool selection and invocation accuracy.
In the remote sensing field, tool decision capabilities enable UAVs to flexibly select and operate onboard tools or attachments based on task requirements, completing specific objectives. Early methods relied heavily on hand-crafted rules and geometry-based strategies, using predefined mission plans for basic tool operations. However, these methods were often scene-specific, lacking generalization and adaptability. With the rise of Deep Reinforcement Learning (DRL), researchers began exploring autonomous learning frameworks to endow agents with tool usage abilities. For example, Levine et al. proposed an end-to-end deep learning framework enabling robots to optimize tool usage strategies based on environmental feedback. Nonetheless, DRL methods are computationally intensive and struggle with task switching or scene variations. Recently, driven by advances in LLMs, a surge of LLM-based agents has emerged. These approaches use LLMs to analyze task instructions and generate specific tool invocation commands. Typical examples include LangChain[108] and AutoGPT[132], which allow agents to select tools dynamically. CLIPort[133] integrates a vision-language model with a tool manipulation controller, enabling robots to execute tool operations efficiently across multiple tasks. Toolformer[134] enhances tool usage capability by training LLMs with self-supervised learning to invoke external APIs and tools. Re-Invoke[135] utilizes the LLM’s query comprehension to extract contextual intent from user inputs and applies multi-perspective similarity ranking to help agents efficiently retrieve and invoke the most relevant tools from large toolsets. Moreover, some approaches aim to embed specialized tools directly into LLMs, enabling integrated task analysis and tool execution. Notable examples include WebGPT[136] and WebCPM[137].
IV Future Work
IV-A Language-Centered Adaptive Alignment of Multimodal Remote Sensing Data
Language-Centered multimodal foundation models leverage the semantic relationships encoded in the language modality to enhance cross-modal interaction, and have achieved rapid progress across various domains. However, the alignment of non-language modalities to language often relies on predefined mapping relationships, typically constructed through manual annotation or automated model generation. This dependency limits the model’s performance in interpretation tasks requiring varied semantic granularity. In human cognition, concept alignment is an adaptive process. For example, once a person learns the concept of ”dog” through language, and subsequently becomes familiar with fine-grained categories like ”Husky,” ”Shiba Inu,” or ”Chihuahua,” they can naturally associate images of specific dogs with the correct fine-grained concepts based on accumulated linguistic knowledge, without requiring a fixed image-to-text mapping for each instance. This illustrates that human cross-modal alignment is not rule-based, but rather enabled by semantic understanding and conceptual association, allowing for generalization and inference across modalities. Modern language-centered multimodal foundation models are also endowed with powerful semantic reasoning capabilities. If such models can harness this reasoning ability to infer possible combinations of visual features from the semantic space, and then match them with the perceived visual inputs, they could achieve adaptive alignment between semantic and visual features. This mechanism would allow the model to extrapolate from limited known alignment examples, dynamically adjust the alignment granularity, and flexibly adapt to the complex requirements of remote sensing interpretation tasks.
IV-B Task Understanding and Knowledge Association under Spatiotemporal Evolution and Dynamic Knowledge Constraints
Remote sensing interpretation tasks are often characterized by significant spatiotemporal dynamics. As observation time progresses and spatial coverage expands, the morphology, distribution, and relationships of geo-objects may undergo complex changes — for instance, urban expansion, vegetation phenological shifts, or dynamic disaster monitoring. Meanwhile, expert knowledge and historical memory are also continuously updated in response to environmental changes or through interaction with users. However, current language-centered approaches to task understanding and knowledge association largely rely on static spatiotemporal assumptions, making them ill-suited for dynamically evolving environments and knowledge bases. This often leads to task misinterpretation and knowledge association breakdown, which in turn compromises the accuracy and generalization capability of interpretation systems. In contrast, the human memory system, particularly through the hippocampal memory mechanism, enables experiential accumulation and iterative knowledge refinement. By continuously integrating new experiences with historical memory, humans enrich their understanding of the world. If such a memory-inspired mechanism could be introduced into large models—allowing them to dynamically update knowledge under spatiotemporal evolution and knowledge constraints—then the adaptability and reasoning ability of remote sensing interpretation models could be significantly enhanced.
IV-C Trustworthy Reasoning and Adaptive Decision-Making for Complex Interpretation Tasks
Currently, MLLMs possess the ability to understand complex interpretation tasks and can explicitly express reasoning processes. However, the reasoning and decision-making processes of existing MLLMs remain relatively isolated from real-world environments. These models primarily rely on internal reasoning chains for self-verification, lacking effective external feedback from the environment. This limitation can result in decisions that are logically valid but infeasible in real-world conditions, which significantly hinders their deployment in low-tolerance domains such as national defense and resource monitoring. This paper argues that intelligent remote sensing interpretation systems must exhibit both trustworthiness and adaptive decision-making capabilities. First, models should be capable of reasoning based on external environmental and task-specific conditions, rather than relying solely on internal recursive verification. Second, they should possess the ability to adaptively optimize decision-making based on execution conditions—such as available tools, environmental constraints, and task-specific preferences—in order to select the most feasible and optimal interpretation path. Therefore, in future research, exploring how to enable MLLMs to reason reliably and make adaptive, executable decisions under complex interpretation tasks and dynamic open environments will be a challenging yet highly promising direction.
IV-D A Language-Centered Interactive System for Autonomous Remote Sensing Image Interpretation
Although language-centered remote sensing interpretation systems are capable of reasoning with world knowledge and logical relationships encoded in language models, their level of intelligence is still constrained by their passive, user-driven interaction mechanisms. Existing systems typically rely on users to explicitly define interpretation tasks, lacking the ability to autonomously explore task demands, identify interpretation challenges, or proactively formulate interpretation strategies in complex scenarios. In contrast, human experts in remote sensing interpretation can actively identify potential interpretation problems based on environmental conditions and accumulated knowledge, then engage in problem-oriented reasoning, hypothesis generation, and solution design. This capability for autonomous discovery of task requirements is currently missing from most language-centered interpretation systems. Therefore, the future of remote sensing interpretation systems is inevitably moving toward the development of remote sensing agents, marking a shift from passive response to autonomous decision-making. Building an agent that can proactively discover interpretation needs, identify problems, and dynamically adjust interpretation strategies based on both environmental inputs and evolving knowledge will be key to further enhancing the intelligence and autonomy of remote sensing interpretation systems.
V Conclusion
Remote sensing image interpretation technology is undergoing a paradigm shift from ”vision-centered perception” to ”language-centered cognition.” This paper traces the dialectical relationship between language and vision in the human cognitive system, revealing the inherent limitations of the current vision-centered paradigm in terms of semantic abstraction, dynamic feedback, and closed-loop cognition. Grounded in the Global Workspace Theory of cognitive science, we propose a novel interpretation paradigm that positions language as the central hub for multimodal representation. This framework not only reconstructs the full perception–cognition–decision loop, but also empowers interpretation systems with cross-modal semantic association and interpretable reasoning through the referential and logical predicate functions of language. Compared to traditional vision-centered approaches, the language-centered framework presents three transformative advantages: The natural language interface for task understanding eliminates barriers in human-machine interaction. The integration of dynamic knowledge graphs overcomes the limitations of perception constrained by ”what is seen is what is known.” The cyclic interaction mechanism of the cognitive workspace enables a qualitative leap from static interpretation to autonomous decision-making. While current technologies still face challenges in multimodal representation alignment and spatiotemporal knowledge evolution modeling, the advancement of a language-centered cyclic interpretation paradigm holds promise for transcending the instrumental limitations of existing systems, steering remote sensing interpretation toward becoming a cognitive agent capable of self-evolution. Future research should focus on deep coupling between language, vision, and physical space, constructing dynamically evolving cognitive schemata under spatiotemporal constraints. Ultimately, this will enable a leap from data-driven to emergent cognitive intelligence in remote sensing, opening a new paradigm of cognitive computing where ”language reconstructs the world.”